OpenStudy (anonymous):

If f(x) = (5x + 3)/(4x-5), x is not equal to 5/4, show that fof is an identity function.

OpenStudy (dumbcow):

you mean like a composite function, f(f(x))

OpenStudy (zarkon):

\[f(x)=\frac{5x + 3}{4x-5}\] \[f(f(x))=\frac{5\left(\frac{5x + 3}{4x-5}\right) + 3}{4\left(\frac{5x + 3}{4x-5}\right)-5}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So will I have to solve it sir

OpenStudy (zarkon):

simplify it

OpenStudy (zarkon):

if done correctly...you will get x

OpenStudy (anonymous):

yes i got x

OpenStudy (zarkon):

good

OpenStudy (anonymous):

SO fof is identity function . Right ?

OpenStudy (zarkon):

not usually

OpenStudy (anonymous):

in this question !

OpenStudy (zarkon):

yes..for this question..but not in general

OpenStudy (anonymous):

ok sir thanks again but please stay online i have to ask 8 more questions

OpenStudy (anonymous):

Please

OpenStudy (zarkon):

there are others here that can also help you :)

OpenStudy (anonymous):

sir please

OpenStudy (anonymous):

will anyone else help me

OpenStudy (dumbcow):

In general \[f(f^{-1}(x)) = x\] this is a special case where \[f^{-1}(x) = f(x)\]

OpenStudy (anonymous):

dumbcow will u help me

OpenStudy (anonymous):

ok

OpenStudy (dumbcow):

post your questions someone will help you